Using the residual income valuation model, compute the price per share based on the following information:
- The book value per share is currently $18.00, and it is estimated to be $19.50 in one year from today and $21.00 in two years from now.
- The consensus of earnings per share for next year is $3.75, and it is $4.25 in two years from now.
- The residual income is forecasted to be zero in the third year and thereafter.
- Based on the CAPM, the required rate of return on the firm is 12%

Question 13 options:

1)

$19.24

2)

$20.94

3)

$21.16

4)

$22.05
Answer #1

Calculating the price of share using  residual income valuation model:

RIRI Vo = Bo +171+r) (1 + r)2

where, B0 = Current Book value = $18

r = Required rate of return = 12%

RI1 = Residual income at year end 1

RI2 = Residual income at year end 2

- Computing Residual income of year end 1 & 2,

RIt = EPSt - (r*Bt-1)

RI1 = 3.75 - (0.12*18) = 3.75 - 2.16 = 1.59

RI2 = 4.25 - (0.12*19.5) = 4.25- 2.34 = 1.91

So, now calculating price per share:

1.59 1.91 Vo = 18 + 1 107(1+0.12)1 *(1 +0.12)2

Vo = 18 + 1.4196 +1.5226

Price per share = $ 20.94

Hence, Option 2

If you need any clarification, you can ask in comments.

If you like my answer, then please up-vote as it will be motivating